OpenStudy (anonymous):

Two identical rubber balls are dropped from different heights. Ball 1 is dropped from a height of 104 feet, and ball 2 is dropped from a height of 294 feet. Write a function for the height of each ball.

OpenStudy (anonymous):

hx = 32.9184 - t * 9.8 m/s/s hy = 89.6112 - t * 9.8 m/s/s t is seconds hx, height of ball 1 hy, height of ball 2

OpenStudy (anonymous):

where did you get those numbers?

OpenStudy (anonymous):

104 and 294 are converted from feet to meters 9.8 m/s/s is the acceleration rate due to gravity

OpenStudy (anonymous):

Considering no air resistance